Ten customers visit C&A Bakery from 8 a.m. to 10 a.m. The customers spend 10, 15, 20, 11, 8, 12, 25, 18, 29, and 32 minutes in the bakery. On average, how many customers are in the bakery from 8 a.m. to 10 a.m.?

   

10

   

1.5

   

90

   

Cannot be determined
Answer #1

Average flow time is (10 + 15 + 20 + 11 + 8 + 12 + 25 + 18 + 29 + 32)/10 = 18 Mins.

8Am- 10Am there are 2 hours

2 hours= 120 min

Average flow rate = 10 customers/120 minutes = 1/12 customer per minute.

Inventory = 18*1/12= 1.5 customers

So Second option is correct.